3000 grams of water(at a temperature of 20 degrees celcius) is mixed with 1000 grams of hot water(at a temperature of 80 degrees celcius) Assuming no heat energy is exchanged with the environment, what is the final equlibrium temperature?
Answer #1

Suppose final temperature is T

Heat gained by 3000 g water = m s (T - 20)

= 3000 × s × (T - 20). -------(1)

Heat lost by 1000 g water = 1000 × s × (80 - T) ------ (2)

Heat lost = heat gained

Equating (1) and (2)

(80 - T ) = 3 ( T - 20)

4 T = 140

T = 35 0C